Portsmouth Company makes upholstered furniture. Its only variable cost is direct materials. The demand for the company's products far exceeds its manufacturing capacity. The bottleneck (or constraint) in the production process is upholstery labor-hours. Information concerning three of Portsmouth's upholstered chairs appears below:

Recliner Sofa Love Seat
  Selling price per unit $ 1,084    $ 1,940    $ 1,325   
  Variable cost per unit $ 700    $ 1,400    $ 900   
  Upholstery labor-hours per unit 6 hours    12 hours    5 hours   

Required:
1.

Portsmouth is considering paying its upholstery laborers hourly compensation, in addition to their usual salaries, to work overtime. Assuming that this extra time would be used to produce sofas, up to how much of an overtime premium per hour should the company be willing to pay to keep the upholstery shop open after normal working hours?

      

2.

A small nearby upholstering company has offered to upholster furniture for Portsmouth at a price of $41 per hour. The management of Portsmouth is confident that this upholstering company’s work is high quality and their craftsmen can work as quickly as Portsmouth’s own craftsmen on the simpler upholstering jobs such as the Love Seat. Should management accept this offer?

Yes
No
Answer #1

SOLUTION

1.

Selling price per unit (A) 1,940
Less: Variable price per unit (B) 1,400
Contribution margin per unit (C=A-B) 540
Labor hours required for one unit (D) 12 hours
Contribution per labor hour (C/D) 45

2. Recliner

Selling price per unit (A) 1,084
Less: Variable price per unit (B) 700
Contribution margin per unit (C=A-B) 384
Labor hours required for one unit (D) 6
Contribution per labor hour (C/D) 64

Love seat

Selling price per unit (A) 1,325
Less: Variable price per unit (B) 900
Contribution margin per unit (C=A-B) 425
Labor hours required for one unit (D) 5
Contribution per labor hour (C/D) 85
Price charged to upholster by another firm 41
Additional margin earned 44

Yes, it should accept the offer.